OpenStudy (anonymous):

My math teacher told me to answer and explain the answer for: 0/0=

OpenStudy (anonymous):

It is "indeterminate", i.e. not defined.

OpenStudy (anonymous):

explain why its not defined

OpenStudy (anonymous):

when there is something over 0 than itll be undefined cuz 8/0 u cant actually divide it

OpenStudy (anonymous):

so any number over 0 is always undefined

OpenStudy (anonymous):

If we think of division as the inverse (the opposite) operation to multiplication, e.g. \[ \frac{6}{2} = 3\]means \[ 2 \times 3 = 6.\]In general, if you have numbers \(a,b,c\), then \[ \frac{a}{b} = c\]means \[ b\times c = a.\]Now, let's say \[ \frac{0}{0} = a.\]Then \(0a = 0\), which allows \(a\) to basically be any number. That's why it's not defined.

OpenStudy (psymon):

By just definition, if you have a division like: (a/b)= c, then this means you can also go backwards and say that c times b = a If you have something like (0/6) = 0, well this is valid because I can go backwards and multiply 0 times 6 to get 0. But if I have the 0 on bottom (6/0) = 1 for example. Well, there is no possibleway for me to go backwards. I cannot say 1 times 0 = 6. I cannot say anything times 0 = 6. This is why it is not possible and therefore undefined.

OpenStudy (psymon):

Looks like I just repeated what was above, lol.

OpenStudy (anonymous):

thanks guys for your help!

OpenStudy (psymon):

As long as it all makes sense xD

OpenStudy (anonymous):

I wish I could give you guys a medal

OpenStudy (psymon):

No biggie. If we're only here for medals then shame on us xD

OpenStudy (anonymous):

I agree with Psymon there :-) @tryder You're welcome!